OpenStudy (anonymous):

How do I solve: 0.008484 / 0.0028 ?

OpenStudy (anonymous):

(8484/1000000)÷(28/10000)= (8484/1000000)*(10000/28)=303/100 = 3.03

OpenStudy (anonymous):

got it :-) Thank you

OpenStudy (anonymous):

yw